For each of the following, find the length of AB
a. A(0,3)B(0,7)
b. A(0,3)B(4,0)

Answer by ewatrrr(24785)


Hi,
a. A(0,3)B(0,7) length 0f AB = 4
b. A(0,3)B(4,0) length of AB = 4^2 + 3^2 = 16+9 =25 = AB^2, Length of AB = 5